OpenStudy (anonymous):

Using DeMoivre's Formula, solve, (1-i)^-8

OpenStudy (anonymous):

You should really try to attempt this one using the techniques covered in the previous question. I'll give you a hint: \(1-i=\sqrt2\left(\frac{\sqrt2}2-i\frac{\sqrt2}2\right)\).

OpenStudy (anonymous):

So it would be like this?